# Hierarchy of timescales in a disordered spin-$1/2$ XX ladder

### v2

@@ Abstract
Understanding the timescales associated with relaxation to equilibrium in closed quantum many-body systems is one of the central focuses in the study of their non-equilibrium dynamics. At late times, these relaxation processes exhibit universal behavior, emerging from the inherent randomness of chaotic Hamiltonians. In this work, we investigate a disordered spin-$1/2$ XX ladder — an experimentally realizable model known for its diffusive dynamics — to explore the connection between transport properties and spectral measures derived solely from the system’s energy levels via these relaxation timescales. We begin by analyzing the spectral form factor, which yields the time when the system begins to follow the random matrix theory (RMT) statistics, known as the RMT time. We then determine the Thouless times — the average times for a local excitation to diffuse across the entire finite system — through the linear-response theory for both spin and energy transport. Our numerical results confirm that the RMT time scales quadratically with system size and upper bounds the Thouless times. Interestingly, we also find that, unlike other non‑integrable models, spin diffusion proceeds faster than energy diffusion.
@@ I Introduction
In the study of non-equilibrium dynamics in closed quantum systems, identifying the relevant timescales is a long-standing problem [1]. Efforts to obtain rigorous bounds on such timescales were reported in [2, 3, 4, 5, 6, 7]. A defining aspect in the non-equilibrium dynamics of closed quantum many-body systems [8] is that relevant relaxation timescales can depend on system size [1, 9]. Most notably, the time to diffuse through a finite system scales quadratically with linear-system size, and is called the Thouless time [10, 11, 12]. Generic non-integrable many-body systems are believed to be described by quantum chaos and eigenstate thermalization hypothesis (ETH) [13, 14, 15, 16, 1]. In this context, the question of how physical timescales are reflected in, e.g., spectral properties [17, 18, 19, 20, 21, 9, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35], survival probabilities [36, 37, 38, 39, 40, 41, 42, 43], the ETH itself [44, 45, 46, 47], or in autocorrelations of local observables [48, 49, 50, 51] has attracted significant attention. Setting these different approaches into a comprehensive context [52] and to establish connections to experiments [53, 54, 55] is an important goal of ongoing research.
@@
To contribute to this objective, we first draw upon quantum chaos [56, 57, 58, 59, 60] — which is closely connected to the ETH through its relationship with random matrix theory (RMT) [1]. One way to study the onset of random matrix behavior in a microscopic model is the utilization of the spectral form factor (SFF) — a widely-used measure to analyze the universal RMT dynamics by examining the spectral fluctuations [57, 58, 61, 59, 62, 63] in quantum many-body physics (see Ref. [64] for a review). This measure does not require the selection of an initial state or an observable as it is solely based on the correlations in the energy spectrum. The onset of this RMT dynamics defines the first timescale that is of particular interest to us — the RMT time. Notably, it has been theoretically established that this timescale exhibits different scaling behaviors with system size [25, 20, 9, 24, 35, 65, 19]. While the RMT time scales quadratically with the system size for a wide range of non-integrable many-body models [25, 21, 20, 9, 22, 24, 32, 35, 65], other scaling behaviors also exist, including other power laws [66], logarithmic [21, 20, 65, 35], or no scaling with system size at all [32, 35]. The broad spectrum of scaling properties can be attributed to the complex and varied mechanisms that govern thermalization processes of the corresponding systems [35, 65, 32, 24, 20, 22]. The RMT time is commonly argued to be identical to the slowest physical relaxation time of the system [64].
@@
Very often, the RMT time is related to diffusion, and hence an $L^{2}$ scaling is expected and indeed observed [25, 20, 9, 22, 24]. For instance, Ref. [67] demonstrates that RMT properties can be utilized to obtain the diffusion constant from level curvature. In our work we aim at a quantitative comparison between the RMT time and the Thouless time, the relevant timescale for diffusion. Transport in interacting one-dimensional models is a long-studied problem (for a review, see, e.g., Refs. [68, 69, 70, 71]). At finite temperatures, generic non-integrable systems are considered to exhibit diffusive dynamics [70] while anomalous transport is found in integrable [72, 73, 70, 74, 75, 76, 77] or weakly-ergodicity breaking systems (see, e.g., Refs. [74, 78]). A straightforward way to compute diffusion constants is the utilization of linear-response theory [70]. How to actually accomplish this in an efficient way in numerical simulations is the topic of ongoing method development [79, 80, 81, 82, 83, 84, 85, 86, 87, 67, 88]. While first experiments with optical lattices investigated transport in the far-from-equilibrium regime [89, 90, 91], recently, direct measurements of diffusion constants in non-integrable models [92] and classifications of transport in integrable models have become possible [75, 77]. In earlier experiments with quantum magnets, both thermal and spin transport of spin chain and ladders systems were studied [93] (see also Ref. [76]).
@@
Crucially, transport phenomena also yield a timescale called the Thouless time — the average time for a local perturbation to spread through a finite system [10, 11, 12]. Its scaling behavior significantly depends on the type of transport: under diffusive dynamics it scales quadratically with the linear-system size, whereas under ballistic dynamics it scales linearly with the system size [94]. (Super)Subdiffusive behavior yields $L^{\gamma}$ with $\gamma>2$ ($1<\gamma<2$) [95].
@@
In the presence of more than one transport channel, there are several Thouless times, e.g., the energy-Thouless time associated with energy transport or the particle-Thouless time associated with particle transport (see, e.g., Ref. [96]). A priori, it is not clear how these Thouless times for energy-, spin- and particle-diffusion relate to each other. In the best studied examples of integrable one-dimensional spin-$1/2$ systems, energy transport is always faster than spin transport, even qualitatively [97, 70].
@@
Figure 1: Sequence of regimes for a diffusive non-integrable many-body quantum spin system with a finite system size $L$. The diffusion of a conserved quantity for such a system is governed by the $L^{2}$-scaling timescale called Thouless time $t_{\mathrm{Th}}$, which represents the average time for a transport quantity to spread through the finite system. Often, energy diffusion is expected to occur more rapidly than spin diffusion, i.e., $t_{\mathrm{Th}}^{(\mathrm{E})}<t_{\mathrm{Th}}^{(\mathrm{S})}$. The random matrix theory (RMT) time $t_{\mathrm{RMT}}$ indicates when the system begins to behave like a random matrix, following universal RMT statistics. This time is expected to be identical (or related) to the longest physical timescale. On top of all of these $L^{2}$-scaling relaxation timescales, the Heisenberg time $t_{\mathrm{H}}$, determined by the mean-level spacing $\overline{\delta E}$, is the longest timescale and exhibits exponential scaling with respect to the system size.
@@
The expected hierarchy of relaxation timescales for a diffusive non-integrable many-body quantum system is illustrated in Fig. 1. The sequence begins with the onset of the hydrodynamic regime, marked by the hydrodynamic-onset time, at which point the system reaches local equilibrium [98, 99]. Perturbations in the density associated with a conserved quantity reach the boundaries of the finite system at their respective Thouless times. After all local perturbations have spread through the finite-dimensional system, the system’s dynamics become featureless and similar to that of random matrices at the RMT time [25, 21, 20, 9, 22, 24, 32, 35, 65, 19, 66]. As a result, the RMT time is expected to be identical to the slowest physical relaxation timescale [64]. Eventually, beyond the Heisenberg time, the system’s dynamics are controlled by the inverse level spacing of the finite-sized system due to the energy-time uncertainty relation, as the system becomes aware of the discrete nature of its energy spectrum [100].
@@
To investigate the hierarchy of these relaxation times, we choose a disordered spin-$1/2$ XX ladder as a study case, as shown in Fig. 2. We induce disorder to overcome the non-self-averaging behavior of the SFF [101]. Transport in the clean case of this model is well-studied, both theoretically [79, 102, 103, 80, 104, 81, 82, 85] and experimentally [92]. Theoretical and experimental results yield comparable results for the diffusion constant [92]. However, these studies are focused exclusively on spin transport, whereas our investigation considers both transport channels in the disordered case. Previous studies of transport looked at a larger family of related spin-ladder models that include XXZ type interactions on the rungs [79], also including disorder that respects a $\mathbb{Z}_{2}$ symmetry associated with exchanging the legs [105].
@@
Our computations reveal that RMT time and Thouless times both scale linearly with $L^{2}$ and therefore, the former appears to be controlled by diffusive hydrodynamics. The RMT time is larger than both Thouless times for energy and spin transport, which we shall discuss further.
@@
Additionally, our results show, for both disordered and clean cases, that the two-leg ladder model possesses a slower energy diffusion than spin diffusion, i.e., the energy-Thouless time is longer than the spin-Thouless time. Other well-studied non-integrable models exhibit the opposite behavior. As an example, we consider a disordered spin-$1/2$ XXZ chain with next-nearest-neighbor couplings [69, 106, 107, 108, 109, 79]. While the RMT time in this model also scales quadratically with system size, here we find the more conventional scenario with a slower spin than energy diffusion. We provide an intuitive explanation for the existence of faster-than-energy spin diffusion in the XX ladder.
This paper is organized as follows: In Sec. II, we begin by introducing the model. Section III describes the methodology used to extract the relaxation timescales from the relevant measures. The results of these measures are then presented in Secs. IV.1 and IV.2, where we discuss the outcomes for varying system sizes of a disordered spin-$1/2$ XX ladder. In Sec. IV.3, we compare the relaxation timescales obtained from these measures for both a disordered spin-$1/2$ XX ladder and, for comparison, a disordered spin-$1/2$ XXZ chain with next-nearest neighbor couplings. Finally, our conclusions are summarized in Sec. V, and additional technical details regarding the computations are provided in the Appendices.
@@ II Model
We consider a system of spin-$1/2$ objects interacting with their nearest-neighbor spins only on the XY plane of the spin space on a two-leg ladder lattice with periodic boundary conditions along the legs, i.e., a spin-$1/2$ XX ladder, as shown in Fig. 2. To make it disordered, we apply random magnetic fields along the Z-direction in spin space, which yields a disorder realization $\boldsymbol{\mathcal{W}}$. Setting $\hbar=1$ throughout the paper, its Hamiltonian can be written as a sum of three different operators as follows:
@@
where $J_{\parallel}$ is the leg-coupling strength, $L$ is the number of the rungs on the lattice, which defines the number of spins, $N=2L$, and $S^{\pm}_{\ell,i}=S^{x}_{\ell,i}\pm iS^{y}_{\ell,i}$ are the spin-$1/2$ ladder operators at the $\ell$-rung and $i$-leg. The second part is responsible for the spin couplings on the rungs, given by
@@
In a clean spin-$1/2$ XX ladder (where $W=0$), the coupling ratio $J_{\perp}/J_{\parallel}$ determines the system’s behavior. When $J_{\perp}/J_{\parallel}=0$, the system consists of two XX chains, which can be mapped to free spinless fermions through the Jordan-Wigner transformation. In the opposite limit where $J_{\perp}/J_{\parallel}\rightarrow\infty$, the system consists of $L$ uncoupled dimers, where each dimer acts as an independent two-spin-$1/2$ system. For any intermediate coupling ratio $J_{\perp}/J_{\parallel}$, the Jordan-Wigner transformation converts the system from spin-1/2 objects into interacting spinless fermions. The model can equivalently be written as a system of hardcore bosons (see, e.g., Refs. [102, 80]).
Moreover, depending on the parameters, the model displays several symmetries. In a system without any disorder $(W=0)$, the model has the symmetries of, for example, the total magnetization, the spin inversion, and lattice translations, and reflection along the legs. Whenever $W\neq 0$, the translation and reflection symmetries are broken because of the introduced random local magnetic fields acting on each spin.
@@ III Relaxation timescales
In this section, we introduce how we extract the relaxation timescales for our quantum many-body system, utilizing measures from different areas of physics. First, we use the spectral form factor to obtain the RMT time. Secondly, we compute diffusion constants to determine the corresponding timescales, referred to as the Thouless times. For each measure mentioned in this section, we describe the numerical methods and technical aspects.
III.1 Random matrix theory time
The RMT time, denoted as $t_{\mathrm{RMT}}$, is a fundamental timescale in the context of quantum chaos, and a widely employed method for extracting this timescale is through the analysis of the spectral form factor. However, despite its popularity, the SFF has a known limitation: it is not self-averaging [120], necessitating the construction of a Hamiltonian ensemble and additional averaging procedures. In our paper, we construct such an ensemble using the disorder realizations.
III.1.1 Spectral form factor (SFF)
Figure 3: Comparison between a typical ensemble-averaged spectral form factor (SFF) of a time-reversal symmetric Hamiltonian (white solid line), $\overline{K}{(\tau)}$, and the SFF of the Gaussian Orthogonal Ensemble (black dashed line), $K_{\mathrm{GOE}}{(\tau)}$, as a function of normalized time $\tau$ at disorder strength $W/J_{\parallel}=1.5$ and coupling ratio $J_{\perp}/J_{\parallel}=1.0$. $\overline{K}{(\tau)}$ exhibits four regimes: quantum, hydrodynamic, universal RMT, and finite size; in contrast, $K_{\mathrm{GOE}}{(\tau)}$ displays only the RMT and finite-size regimes [58, 59, 60].
As shown in Fig. 3, a typical SFF of a generic model generally exhibits four distinct regimes over time: quantum, hydrodynamic, universal RMT, and finite size. Initially, in the non-universal quantum-dynamic regime, it starts at a finite value and smoothly decreases. Then, this is followed by the hydrodynamic regime. Upon completion of these two initial regimes, the SFF transitions into the RMT-dynamic regime. This regime is characterized by a universal behavior, as it begins to follow the SFF of an eigenvalue distribution from RMT (see Ref. [17] and references therein). The time at which this transition occurs is referred to as the normalized RMT time, $\tau_{\mathrm{RMT}}$, marking the onset of universal RMT behavior. Finally, at the end of this ramp, it saturates to a constant value at the normalized Heisenberg time, $\tau_{\mathrm{H}}$. It is inversely proportional to the mean-level spacing and is located at $2\pi$ on the time axis due to the spectral unfolding procedure.

### v3

@@ Abstract
Understanding the timescales associated with relaxation to equilibrium in closed quantum many-body systems is one of the central focuses in the study of their nonequilibrium dynamics. At late times, these relaxation processes exhibit universal behavior, emerging from the inherent randomness of chaotic Hamiltonians. In this work, we investigate a disordered spin-$1/2$ XX ladder—an experimentally realizable model known for its diffusive dynamics—to explore the connection between transport properties and spectral measures derived solely from the energy levels of the system via these relaxation timescales. We begin by analyzing the spectral form factor, which yields the time when the system begins to follow the random matrix theory (RMT) statistics, known as the RMT time. We then determine the Thouless times—the average times for a local excitation to diffuse across the entire finite system—through the linear-response theory for both spin and energy transport. Our numerical results confirm that the RMT time scales quadratically with system size and upper bounds the Thouless times. Interestingly, we also find that, unlike other nonintegrable models, spin diffusion proceeds faster than energy diffusion.
@@ I Introduction
In the study of nonequilibrium dynamics in closed quantum systems, identifying the relevant timescales is a long-standing problem [1]. Efforts to obtain rigorous bounds on such timescales were reported in Refs. [2, 3, 4, 5, 6, 7]. A defining aspect in the nonequilibrium dynamics of closed quantum many-body systems [8] is that relevant relaxation timescales can depend on system size [1, 9]. Most notably, the time to diffuse through a finite system scales quadratically with linear system size and is called the Thouless time [10, 11, 12]. Generic nonintegrable many-body systems are believed to be described by quantum chaos and the eigenstate thermalization hypothesis (ETH) [13, 14, 15, 16, 1]. In this context, the question of how physical timescales are reflected in, e.g., spectral properties [17, 18, 19, 20, 21, 9, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35], survival probabilities [36, 37, 38, 39, 40, 41, 42, 43], the ETH itself [44, 45, 46, 47], or in autocorrelations of local observables [47, 48, 49, 50] has attracted significant attention. Setting these different approaches into a comprehensive context [51] and establishing connections to experiments [52, 53, 54] are important goals of ongoing research.
@@
To contribute to this objective, we first draw upon quantum chaos [55, 56, 57, 58, 59]—which is closely connected to the ETH through its relationship with random matrix theory (RMT) [1]. One way to study the onset of random matrix behavior in a microscopic model is the utilization of the spectral form factor (SFF)—a widely-used measure to analyze the universal RMT dynamics by examining the spectral fluctuations [56, 57, 60, 58, 61, 62] in quantum many-body physics (see Ref. [63] for a review). This measure does not require the selection of an initial state or an observable as it is solely based on the correlations in the energy spectrum. The onset of this RMT dynamics defines the first timescale that is of particular interest to us—the RMT time. Notably, it has been theoretically established that this timescale exhibits different scaling behaviors with system size [25, 20, 9, 24, 35, 64, 19]. While the RMT time scales quadratically with the system size for a wide range of nonintegrable many-body models [25, 21, 20, 9, 22, 24, 32, 35, 64], other scaling behaviors also exist, including other power laws [65], logarithmic [21, 20, 64, 35], or no scaling with system size at all [32, 35]. The broad spectrum of scaling properties can be attributed to the complex and varied mechanisms that govern thermalization processes of the corresponding systems [35, 64, 32, 24, 20, 22]. The RMT time is commonly argued to be identical to the slowest physical relaxation time of the system [63].
@@
Very often, the RMT time is related to diffusion, and hence an $L^{2}$ scaling is expected and indeed observed [25, 20, 9, 22, 24] ($L$ is the linear system size). For instance, in Ref. [66], RMT properties are utilized to obtain the diffusion constant from level curvature. In our work we aim at a quantitative comparison between the RMT time and the Thouless time, the relevant timescale for diffusion. Transport in interacting one-dimensional models is a long-studied problem (for a review, see, e.g., Refs. [67, 68, 69, 70]). At finite temperatures, generic nonintegrable systems are considered to exhibit diffusive dynamics [69] while anomalous transport is found in integrable [71, 72, 69, 73, 74, 75, 76] or weakly ergodicity-breaking systems (see, e.g., Refs. [73, 77]). A straightforward way to compute diffusion constants is the utilization of linear-response theory [69]. How to actually accomplish this in an efficient way in numerical simulations is the topic of ongoing method development [78, 79, 80, 81, 82, 83, 84, 85, 86, 66, 87]. While experiments with optical lattices investigated transport in the far-from-equilibrium regime [88, 89, 90], recently, direct measurements of diffusion constants in nonintegrable models [91] and classifications of transport in integrable models have become possible [74, 76]. In earlier experiments with quantum magnets, both thermal and spin transport of spin chain and ladders systems were studied [92] (see also Ref. [75]).
@@
Crucially, transport phenomena also yield a timescale called the Thouless time—the average time for a local perturbation to spread through a finite system [10, 11, 12]. Its scaling behavior significantly depends on the type of transport: Under diffusive dynamics it scales quadratically with the linear system size, whereas under ballistic dynamics it scales linearly with the system size [93]. (Super)Subdiffusive behavior yields $L^{\gamma}$ with $\gamma>2$ ($1<\gamma<2$) [94].
@@
In the presence of more than one transport channel, there are several Thouless times, e.g., the energy-Thouless time associated with energy transport or the particle-Thouless time associated with particle transport (see, e.g., Ref. [95]). A priori, it is not clear how these Thouless times for energy, spin and particle diffusion relate to each other. In the best-studied examples of integrable one-dimensional spin-$1/2$ systems, energy transport is always faster than spin transport, even qualitatively [96, 69].
@@
Figure 1: Sequence of regimes for a diffusive nonintegrable many-body quantum spin system with a finite system size $L$. The diffusion of a conserved quantity for such a system is governed by the $L^{2}$-scaling timescale called Thouless time $t_{\mathrm{Th}}$, which represents the average time for a transport quantity to spread through the finite system. Often, energy diffusion is expected to occur more rapidly than spin diffusion, i.e., $t_{\mathrm{Th}}^{(\mathrm{E})}<t_{\mathrm{Th}}^{(\mathrm{S})}$. The RMT time $t_{\mathrm{RMT}}$ indicates when the system begins to behave like a random matrix, following universal RMT statistics. This time is expected to be identical (or related) to the longest physical timescale. On top of all of these $L^{2}$-scaling relaxation timescales, the Heisenberg time $t_{\mathrm{H}}$, determined by the mean-level spacing $\overline{\delta E}$, is the longest timescale and exhibits exponential scaling with respect to the system size.
@@
The expected hierarchy of relaxation timescales for a diffusive nonintegrable many-body quantum system is illustrated in Fig. 1. The sequence begins with the onset of the hydrodynamic regime, marked by the hydrodynamic-onset time, at which point the system reaches local equilibrium [97, 98]. Perturbations in the density associated with a conserved quantity reach the boundaries of the finite system at their respective Thouless times. After all local perturbations have spread through the finite-dimensional system, the dynamics of the system become featureless and similar to that of random matrices at the RMT time [25, 21, 20, 9, 22, 24, 32, 35, 64, 19, 65]. As a result, the RMT time is expected to be identical to the slowest physical relaxation timescale [63]. Eventually, beyond the Heisenberg time, the dynamics of the system are controlled by the inverse level spacing of the finite-sized system due to the energy-time uncertainty relation, as the system becomes aware of the discrete nature of its energy spectrum [99].
@@
To investigate the hierarchy of these relaxation times, we choose a disordered spin-$1/2$ XX ladder as a study case, as shown in Fig. 2. We induce disorder to overcome the non-self-averaging behavior of the SFF [100]. Transport in the clean case of this model is well studied, both theoretically [78, 101, 102, 79, 103, 80, 81, 84] and experimentally [91]. Theoretical and experimental results yield comparable results for the diffusion constant [91]. However, these studies are focused exclusively on spin transport, whereas in our investigation, we consider both transport channels in the disordered case. Authors of previous studies of transport looked at a larger family of related spin-ladder models that include XXZ-type interactions on the rungs [78], also including disorder that respects $\mathbb{Z}_{2}$ symmetry associated with exchanging the legs [104].
@@
Our computations reveal that RMT and Thouless times both scale linearly with $L^{2}$ and therefore, the former appears to be controlled by diffusive hydrodynamics. The RMT time is larger than both Thouless times for energy and spin transport, which we shall discuss further.
@@
Additionally, our results show, for both disordered and clean cases, that the two-leg ladder model possesses a slower energy diffusion than spin diffusion, i.e., the energy-Thouless time is longer than the spin-Thouless time. Other well-studied nonintegrable models exhibit the opposite behavior [105, 106, 107]. As an example, we consider a disordered spin-$1/2$ XXZ chain with next-nearest-neighbor couplings [68, 108, 109, 110, 111, 78]. While the RMT time in this model also scales quadratically with system size, here we find the more conventional scenario with a slower spin than energy diffusion. We provide an intuitive explanation for the existence of faster-than-energy spin diffusion in the XX ladder.
This paper is organized as follows: In Sec. II, we begin by introducing the model. Sec. III describes the methodology used to extract the relaxation timescales from the relevant measures. The results of these measures are then presented in Secs. IV.1 and IV.2, where we discuss the outcomes for varying system sizes of a disordered spin-$1/2$ XX ladder. In Sec. IV.3, we compare the relaxation timescales obtained from these measures for both a disordered spin-$1/2$ XX ladder and, for comparison, a disordered spin-$1/2$ XXZ chain with next-nearest-neighbor couplings. Finally, our conclusions are summarized in Sec. V, and additional technical details regarding the computations are provided in the Appendixes.
@@ II Model
We consider a system of spin-$1/2$ objects interacting with their nearest-neighbor spins only on the XY plane of the spin space on a two-leg ladder lattice with periodic boundary conditions along the legs, i.e., a spin-$1/2$ XX ladder, as shown in Fig. 2. To make it disordered, we apply random magnetic fields along the z-direction in spin space, which yields a disorder realization $\boldsymbol{\mathcal{W}}$. For the rest of the paper, we mostly avoid using $\boldsymbol{\mathcal{W}}$ for ease of complex notation. Setting $\hbar=1$ throughout the paper, its Hamiltonian can be written as a sum of three different operators as follows:
@@
where $J_{\parallel}$ is the leg-coupling strength, $L$ is the number of the rungs on the lattice, which defines the number of spins $N=2L$, and $S^{\pm}_{\ell,i}=S^{x}_{\ell,i}\pm iS^{y}_{\ell,i}$ are the spin-$1/2$ ladder operators at $\ell$ rung and $i$ leg. The second part is responsible for the spin couplings on the rungs, given by
@@
In a clean spin-$1/2$ XX ladder (where $W=0$), the coupling ratio $J_{\perp}/J_{\parallel}$ determines the behavior of the system. When $J_{\perp}/J_{\parallel}=0$, the system consists of two XX chains, which can be mapped to free spinless fermions through the Jordan-Wigner transformation. In the opposite limit where $J_{\perp}/J_{\parallel}\rightarrow\infty$, the system consists of $L$ uncoupled dimers, where each dimer acts as an independent two-spin-$1/2$ system. For any intermediate coupling ratio $J_{\perp}/J_{\parallel}$, the Jordan-Wigner transformation converts the system from spin-1/2 objects into interacting spinless fermions. The model can equivalently be written as a system of hardcore bosons (see, e.g., Refs. [101, 79]).
Moreover, depending on the parameters, the model displays several symmetries. In a system without any disorder $(W=0)$, the model has the symmetries of, for example, the total magnetization, the spin inversion, lattice translations, and reflection along the legs. Whenever $W\neq 0$, the translation and reflection symmetries are broken because of the introduced random local magnetic fields acting on each spin.
@@ III Relaxation timescales
In this section, we introduce how we extract the relaxation timescales for our quantum many-body system, utilizing measures from different areas of physics. First, we use the SFF to obtain the RMT time. Secondly, we compute diffusion constants to determine the corresponding timescales, referred to as the Thouless times. For each measure mentioned in this section, we describe the numerical methods and technical aspects.
III.1 RMT time
The RMT time, denoted as $t_{\mathrm{RMT}}$, is a fundamental timescale in the context of quantum chaos, and a widely employed method for extracting this timescale is through the analysis of the SFF. However, despite its popularity, the SFF has a known limitation: It is not self-averaging [122], necessitating the construction of a Hamiltonian ensemble and additional averaging procedures. In our paper, we construct such an ensemble using the disorder realizations.
III.1.1 SFF
Figure 3: Comparison between a typical ensemble-averaged SFF of a time-reversal symmetric Hamiltonian (white solid line) $\overline{K}{(\tau)}$ and the SFF of the Gaussian orthogonal ensemble (black dashed line) $K_{\mathrm{GOE}}{(\tau)}$ as a function of normalized time $\tau$ at disorder strength $W/J_{\parallel}=1.5$ and coupling ratio $J_{\perp}/J_{\parallel}=1.0$. $\overline{K}{(\tau)}$ exhibits four regimes: quantum, hydrodynamic, universal RMT, and finite size; in contrast, $K_{\mathrm{GOE}}{(\tau)}$ displays only the RMT and finite-size regimes [57, 58, 59].
As shown in Fig. 3, a typical SFF of a generic model generally exhibits four distinct regimes over time: quantum, hydrodynamic, universal RMT, and finite size. Initially, in the nonuniversal quantum regime, it starts at a finite value and smoothly decreases. Then this is followed by the hydrodynamic regime. Upon completion of these two initial regimes, the SFF transitions into the RMT regime. This regime is characterized by universal behavior, as it begins to follow the SFF of an eigenvalue distribution from RMT (see Ref. [17] and references therein). The time at which this transition occurs is referred to as the normalized RMT time $\tau_{\mathrm{RMT}}$, marking the onset of universal RMT behavior. Finally, at the end of this ramp, it saturates to a constant value at the normalized Heisenberg time $\tau_{\mathrm{H}}$. It is inversely proportional to the mean-level spacing and is located at $2\pi$ on the time axis due to the spectral unfolding procedure.